OpenStudy (sharkfinsoup):

Assume f(x) = 2x + 8 and g(x) = 5. What is the value of (g o f)(3)?

OpenStudy (sharkfinsoup):

This is the process I used and answer I got. Is it correct? (g o f) (x)= g(f(x)) 5(2x+8) 10x+40 3(10x+40) 30x+120

OpenStudy (unklerhaukus):

g(x) is not equal to 5x g(x) is equal to 5

OpenStudy (sharkfinsoup):

So I can't multiply 2x+8 by 5?

OpenStudy (unklerhaukus):

g(x) = 5 g(2x+8) = 5 g(2(3)+8) = 5

OpenStudy (sharkfinsoup):

so it would be g(14)=5? Where do I go from there?

OpenStudy (unklerhaukus):

(g o f)(3) = g(f(3)) = g(2(3)+8) = g(14) = 5

OpenStudy (sharkfinsoup):

so 5 is the final answer?

OpenStudy (unklerhaukus):

yes

OpenStudy (sharkfinsoup):

I understand that g(x) = 5 but how do you go from g(14) to 5? What's the point of the calculation if the answer given to you in the question?

OpenStudy (unklerhaukus):

The function g, takes an input (say x) and returns the number 5 every time. This kind of question is useful if you are not use to the function(input) notation. You might like to compare this result (g o f)(3), with the result of (f o g)(3) and notice that they are quite different results.

OpenStudy (sharkfinsoup):

Ok that makes sense. Thank you so much